What is the vapor pressure in mm


 Hg of a solution prepared by dissolving 25.0 g


 of ethyl alcohol (C

2

H

5

OH)






 in 100.0 g


 of water at 25

C?



The vapor pressure of pure

water is 23.8mmHg,



 and the vapor pressure of ethyl alcohol is 61.2 mmHg




at 25

C.




(b) What is the vapor pressure of the solution if 25.0 g


 of water is dissolved in 100.0 g


 of ethyl alcohol at 25

C?



Answer




1
Expert's answer
2021-12-21T14:44:02-0500


(a)


Molar mass of water "=18.015"

Molar mass of Ethyl alcohol"=46.07"


Moles of water "=\frac{100}{18.015}=5.551"


Moles of "C_2H_5OH=\frac{25}{46.07}=0.543"


Total moles "=6.094moles"


"P_{water}=\frac{5.551}{6.094}×23.8=21.679"


"P_{alcohol}=\frac{0.543}{6.094}×61.2=5.453"


Vapor pressure of solution "=21.679+5.453"

"=27.132mmHg"



(b)


Moles of water "=\frac{25}{18.015}=1.388"


Moles of ethyl alcohol"=\frac{100}{46.07}=2.171"


Total moles "=3.559"


"P_{water}=\frac{1.388}{3.559}×23.8=9.282"


"P_{Alcohol}=\frac{2.171}{3.559}×61.2=37.332"

Vapor pressure of solution

"=9.282+37.332"

"=46.614mmHg"